# SBoRA: Low-Rank Adaptation with Regional Weight Updates

### v2

@@ Abstract
This paper introduces Standard Basis LoRA (SBoRA), a novel parameter-efficient fine-tuning approach for Large Language Models that builds upon the pioneering works of Low-Rank Adaptation (LoRA) and Orthogonal Adaptation. SBoRA further reduces the computational and memory requirements of LoRA while enhancing learning performance. By leveraging orthogonal standard basis vectors to initialize one of the low-rank matrices, either $\mathbf{A}$ or $\mathbf{B}$, SBoRA enables regional weight updates and memory-efficient fine-tuning. This approach gives rise to two variants, SBoRA-FA and SBoRA-FB, where only one of the matrices is updated, resulting in a sparse update matrix $\mathrm{\Delta}\mathbf{W}$ with a majority of zero rows or columns. Consequently, the majority of the fine-tuned model’s weights $(\mathbf{W}_{0}+\mathrm{\Delta}\mathbf{W})$ remain unchanged from the pre-trained weights. This characteristic of SBoRA, wherein regional weight updates occur, is reminiscent of the modular organization of the human brain, which efficiently adapts to new tasks. Our empirical results demonstrate the superiority of SBoRA-FA over LoRA in various fine-tuning tasks, including commonsense reasoning and arithmetic reasoning. Furthermore, we evaluate the effectiveness of QSBoRA on quantized LLaMA models of varying scales, highlighting its potential for efficient adaptation to new tasks. Code is available at https://github.com/cityuhkai/SBoRA
@@ 1 Introduction
Large language models (LLMs) and diffusion models have become essential components of natural language processing [1, 25] and multimodal AI applications [14, 16]. Full fine-tuning (FFT) has been shown to significantly improve their performance in various downstream tasks [28] or introduce new concepts. However, fine-tuning pre-trained models, especially LLMs, by updating all parameters is computationally costly. As depicted in Fig. 1(a), FFT involves directly updating the high-dimensional pre-trained weight matrix $\mathbf{W}_{0}$.
To address this issue, parameter-efficient fine-tuning (PEFT) methods [9] have gained momentum, focusing on updating only a small fraction of parameters, like adapter [7, 9, 19], prompt tuning [13, 21, 26]. Among these methods, Low-Rank Adaptation (LoRA) has emerged as a groundbreaking technique for LLMs. LoRA [10] introduces a parallel low-rank adapter to the weights of linear layers as shown in Fig. 1(b), reducing memory overhead and computational costs during fine-tuning. However, there are still limitations to LoRA, including activation memory consumption and a performance gap compared to FFT. These limitations have led to the development of many LoRA variants [2, 3, 4, 6, 17, 22, 23, 27, 29, 31, 32].
@@
In this work, we introduce Standard Basis LoRA (SBoRA), a novel approach that further reduces the computational and memory requirements while enhancing the learning capacity. SBoRA achieves this by selectively updating specific rows or columns of the pre-trained weight matrix $\mathbf{W}_{0}$, preserving most of the base model’s weights. The resulting update matrix $\mathrm{\Delta}\mathbf{W}=\mathbf{BA}$ mainly consists of zero rows or columns, indicating that most of the fine-tuned model’s weights $(\mathbf{W}^{\prime}=\mathbf{W}_{0}+\mathrm{\Delta}\mathbf{W})$ remain identical to the pre-trained weights. This localized learning process parallels the modular organization of the human brain, where specific cognitive functions are localized in distinct regions [5]. For example, the hippocampus is responsible for episodic memory [30] while the prefrontal cortex is crucial for working memory and decision-making [12]. By mirroring the modular functionality of the human brain, SBoRA is expected to improve knowledge retention and facilitate efficient adaptation.
Let’s assume that the pre-trained weight matrix $\mathbf{W}_{0}$, the projection-down matrix $\mathbf{A}$, and the projection-up matrix $\mathbf{B}$ are given. SBoRA adopts a unique approach, utilizing orthogonal standard basis vectors (one-hot vectors) to construct its projection matrices. These matrices, designated as $\mathbf{A}_{sb}$ for SBoRA-FA and $\mathbf{B}_{sb}$ for SBoRA-FB, are illustrated in Figures 1(c) and 1(d), respectively. This efficiency stems from the elimination of the need to store coefficients for one of the projection matrices, resulting in approximately 50% reduction in memory usage. In SBoRA-FA, besides the frozen weight matrix $\mathbf{W}_{0}$, the $\mathbf{A}_{sb}$ is fixed and only updates $\mathbf{B}$ during fine-tuning. The weight changes ($\mathrm{\Delta}\mathbf{W}=\mathbf{BA}_{sb}$) reside in a low-rank subspace spanned by $\mathbf{A}_{sb}$, and the projection-down transformation can be efficiently achieved by selecting a subset of input vector samples. SBoRA-FB operates similarly, using a pre-defined standard basis projection-up matrix $\mathbf{B}_{sb}$ and updating only matrix $\mathbf{A}$. Both SBoRA-FA/FB align with LoRA, reducing computational overhead during fine-tuning and introducing no extra inference latency.
@@
The summary of our contributions is as follows:
  • Introduces regional weight updates in PEFT that mirroring the human brain’s ability to store specific skills in distinct, specialized regions.
[…]
  • Introduces sample-based multiplication to realize the efficient SBoRA-FA projection-down and SBoRA-FB projection-up transformations that reduce computational complexity by eliminating one matrix multiplication operation.
@@ 2 Related Work
In deep learning, fine-tuning is a fundamental concept that enables leveraging pre-trained models as a starting point for new tasks [7, 9]. This approach is based on the idea that a model trained on one task can be adapted to perform well on another related task.
2.1 Full Fine-Tuning (FFT)
FFT involves updating all of the parameters of a pre-trained model on a new dataset for a specific task as shown in Fig. 1(a). This process adjusts the model’s weights to fit the new task, leading to improved performance. However, the FFT approach has several limitations. Notably, it necessitates storing and deploying numerous copies of large models tailored to different tasks, which can be prohibitively storage-intensive. Furthermore, FFT is computationally costly, particularly for LLMs, and can be overly time-consuming.
2.2 Parameter-Efficient Fine-Tuning (PEFT)
To address the limitations of FFT, researchers have proposed PEFT techniques [9]. These methods adapt only a small subset of a model’s parameters to a new task, aiming to achieve comparable performance to FFT while reducing computational and memory requirements. Early approaches include the use of adapter layers, which add small, trainable modules to each Transformer layer. Another approach is prefix tuning [15], which optimizes a few continuous "prefix" vectors prepended to the input sequence. Both of these approaches significantly reduce the memory and computation needed to fine-tune large models. However, they also have their limitations, such as extra inference latency in the case of adapter layers, and difficulty in optimization for prefix tuning.
@@ 2.3 Low-Rank Adaptation (LoRA)
LoRA [10], a prominent PEFT method, has garnered significant attention due to its effectiveness and efficiency. As depicted in Fig. 1(b), the fundamental concept underlying LoRA involves approximating full-rank updates with low-rank updates within the FFT domain. Specifically, given a pre-trained parameter matrix $\mathbf{W}_{0}\in\mathbb{R}^{d\times k}$, LoRA utilizes two low-rank matrices, $\mathbf{A}\in\mathbb{R}^{r\times k}$ and $\mathbf{B}\in\mathbb{R}^{d\times r}$, to compute the weight update $\mathrm{\Delta}\mathbf{W}\in\mathbb{R}^{d\times k}$. The output vector $\mathbf{h}\in\mathbb{R}^{d\times 1}$ of a LoRA linear layer with input $\mathbf{x}\in\mathbb{R}^{k\times 1}$ can be expressed as
@@
LoRA ensures $\mathrm{\Delta}\mathbf{W}$ is initialized to zero by initializing $\mathbf{A}$ with a uniform distribution and $\mathbf{B}$ with zero. The low-rank decomposition of $\mathrm{\Delta}\mathbf{W}$ into $\mathbf{BA}$ indicates that the rank of $\mathrm{\Delta}\mathbf{W}$ is at most $r$, significantly lower than full-rank updating in FFT, with $r$ being much smaller than the minimum of $d$ and $k$.
LoRA’s low-rank updating approach has demonstrated comparable performance to full-rank updating in tasks like text classification and instruction tuning. It reduces memory usage and computational requirements by decreasing the number of trainable parameters needed. LoRA injects low-rank matrices directly into existing layers without adding extra layers, resulting in no additional inference latency. Switching out different LoRA modules is easy, simplifying deployment in multi-task scenarios. Empirical studies show that LoRA can match or exceed FFT performance using only a fraction of the total parameters, reducing the need for high-performance hardware.
@@ 3.2 Regional Weight Update
An interesting property of the proposed SBoRA is that the merged weight matrix $\mathbf{W}^{\prime}$ of the fine-tuned model is only regionally updated, with most of the weights remaining unchanged from the original pre-trained weight $\mathbf{W}_{0}$ as depicted in Fig. 2. Specifically, the fine-tuned weight $\mathbf{W}^{\prime}$ of SBoRA-FA can be represented as:
@@ 4 Complexity Analysis of SBoRA
The implementation of SBoRA offers a high degree of flexibility, and we propose sampling-based multiplication, sampleMul, to achieve lower memory and computational requirements. Essentially, both SBoRA-FA/FB can be implemented similarly to conventional LoRA through matrix multiplication, with just frozen the standard basis matrix $\mathbf{A}_{sb}$/$\mathbf{B}_{sb}$. As a result, the computational complexity and memory usage of SBoRAs are similar to those of LoRA. However, we can optimize memory usage by representing the standard basis matrix $\mathbf{A}_{sb}$/$\mathbf{B}_{sb}$ as a $r$-dimensional vector with integer coefficients corresponding to the indices of the selected subspace basis vectors. The basis vector indices are initialized by randomly sampling $r$ unique numbers between 1 and $k$ for SBoRA-FA or between 1 and $d$ for SBoRA-FB. This implementation of SBoRA achieves a significant reduction in parameter storage, approximately 50%, while also saving around half of the gradient storage due to the standard basis matrix $\mathbf{A}_{sb}$/$\mathbf{B}_{sb}$ being frozen. Table 1 provides an analysis of the memory requirements for SBoRA compared to LoRA.
@@
To reduce computational complexity, we can leverage one-hot standard basis vectors to simplify the projection transformation, effectively converting it into a component sampling process that eliminates the need for multiplication operations. Specifically, we select components of the input vector based on the indices of the standard basis vectors. This allows us to sample the input vector into the projection-down vector based on the corresponding indices in $\mathbf{A}_{sb}$. The forward process of SBoRA-FA with sampleMul can be expressed as:
@@
where $\mathbf{x}[seq,\mathbf{A}_{sb}]$ is a component sampling process method for implementing $\mathbf{A}_{sb}\mathbf{x}$. This approach is also applied to the case of SBoRA-FB in the projection-up transformation. We can utilize the $index\_add$($\cdot$) operation to incorporate the non-zero values from the output projection-up vector into the pre-trained weight output vector ($\mathbf{W}_{0}\mathbf{x}$) based on the corresponding indices in $\mathbf{B}_{sb}$. The forward process of SBoRA-FB with sampleMul can be expressed as:
@@
Notably, for SBoRA-FB, projection-down vector $\mathbf{z}$ is directly added into the pre-trained weight output vector through index addition method, eliminating the need for matrix multiplication operations. During this process, two matrix multiplication operations $\mathbf{B}(\mathbf{A}_{sb}\mathbf{x})$ / $\mathbf{Ax}$ and $\mathbf{W}_{0}\mathbf{x}$, and one matrix addition operation of $\mathbf{W}_{0}\mathbf{x}+\mathbf{BA}_{sb}\mathbf{x}$ / $\mathbf{W}_{0}\mathbf{x}+\mathbf{B}_{sb}\mathbf{Ax}$ are involved.
It reduces computation complexity by saving one matrix multiplication operation compared to LoRA[10]. For our experiments, we used the SBoRA sampleMul implementation due to its lower computational complexity and memory usage compared to traditional LoRA methods. Table 2 provides an analysis of computational requirements for SBoRA using sampleMul compared to LoRA, assuming input $\mathbf{x}$ in the shape of $[d\times 1]$.
Table 2: Computational analysis and comparison of LoRA and SBoRA’s sampleMul implementation, with a pre-trained weight shape of $k\times d$ and a LoRA rank of $r$. Notably, in the SBoRA_sampleMul implementations, the number of multiplication and addition operations is significantly reduced.
@@ 5 Experiment
We conducted a comprehensive evaluation of SBoRA’s effectiveness across a range of tasks. To begin with, we compared the performance of SBoRA-FA and SBoRA-FB to that of several other well-known PEFT methods by fine-tuning LLMs, specifically LLaMA-7B, on both commonsense reasoning tasks and arithmetic reasoning tasks. Additionally, given the recent release of LLaMA3 [18], we expanded our evaluation to assess SBoRA’s performance using LLaMA3-8B. Furthermore, recognizing the growing interest in quantized models, we also evaluated SBoRA on the quantized versions of LLaMA-7B LLaMA-13B and LLaMA3-8B models, utilizing the techniques introduced in QLoRA [4].
@@ 5.1 Evaluating SBoRA on Commonsense Reasoning Tasks
We evaluated SBoRA-FA/FB against LoRA, and DoRA approach on the LLaMA-7B and LLaMA3-8B language models for commonsense reasoning tasks. We followed the evaluation method introduced in DoRA [17] and LLM-Adapter [11] which used a comprehensive training dataset created by aggregating eight tasks, evaluating models on individual testing datasets for each task. In addition, the results of GPT-3.5 are included for reference as shown in Table 3.
For a comprehensive comparison, we initially fine-tuned models with SBoRA-FA and SBoRA-FB, both using a double rank of 64 to maintain a nearly identical number of trainable parameters (TP) as LoRA (rank=32), with only additional storage for standard basis indices. We then conducted experiments with rank 32 for SBoRA-FA/FB and 64 for other baselines to investigate if SBoRA could achieve better performance with half the trainable parameters. Consistency was maintained by using one training epoch and learning rate of 2e-4 for all experiments.
In LLaMA-7B, SBoRA-FA (rank=64) outperforms all baseline methods with a similar number of trainable parameters, including LoRA (rank=32), and DoRA (rank=32). SBoRA-FA (rank=32) achieves commendable performance, matching the accuracy of DoRA (rank=32) and trailing behind the best accuracy of LoRA (rank=32) by only 1.1%. Despite having approximately half the trainable parameters, SBoRA performs comparably to models with the same rank. While in LLaMA3-8B, the top two results are achieved by SBoRA-FA with rank 64 and 32, respectively.
Importantly, under identical settings and rank with NVIDIA Geforce RTX 4090, SBoRA-FA/FB significantly reduces training time by half compared to DoRA. Fine-tuning with DoRA takes around 7-8 hours, while achieving comparable or better performance with SBoRA can be done in just 3-4 hours. The notable computational efficiency of SBoRA saves a substantial amount of time.
Table 3: Comparison of LLaMA-7B and LLaMA3-8B with different PEFT methods, and evaluating on the commonsense reasoning task. The first row list the results of GPT-3.5 for reference. We report the accuracy (%) results for each of the eight sub-tasks as well as the average accuracy (%), higher is better for all metrics. The column headers indicates TP for the number of trainable parameters and $r$ for the rank.

### v3

@@ Abstract
This paper introduces Standard Basis LoRA (SBoRA), a novel parameter-efficient fine-tuning approach for Large Language Models that builds upon the pioneering works of Low-Rank Adaptation (LoRA) and Orthogonal Adaptation. SBoRA reduces the number of trainable parameters by half or doubles the rank with the similar number of trainable parameters as LoRA, while improving learning performance. By utilizing orthogonal standard basis vectors to initialize one of the low-rank matrices (either $\mathbf{A}$ or $\mathbf{B}$), SBoRA facilitates regional weight updates and memory-efficient fine-tuning. This results in two variants, SBoRA-FA and SBoRA-FB, where only one of the matrices is updated, leading to a sparse update matrix $\mathrm{\Delta}\mathbf{W}$ with predominantly zero rows or columns. Consequently, most of the fine-tuned model’s weights $(\mathbf{W}_{0}+\mathrm{\Delta}\mathbf{W})$ remain unchanged from the pre-trained weights, akin to the modular organization of the human brain, which efficiently adapts to new tasks. Our empirical results demonstrate the superiority of SBoRA-FA over LoRA in various fine-tuning tasks, including commonsense reasoning and arithmetic reasoning. Furthermore, we evaluate the effectiveness of QSBoRA on quantized LLaMA models of varying scales, highlighting its potential for efficient adaptation to new tasks. Code is available at https://github.com/cityuhkai/SBoRA
@@ 1 Introduction
Large language models (LLMs) and diffusion models have become essential components of natural language processing [1, 25] and multimodal AI applications [14, 16]. Full fine-tuning (FFT) has been shown to significantly improve their performance in various downstream tasks [12, 28] or introduce new concepts. However, fine-tuning pre-trained models, especially LLMs, by updating all parameters is computationally costly. As depicted in Fig. 1(a), FFT involves directly updating the high-dimensional pre-trained weight matrix $\mathbf{W}_{0}$.
To mitigate this challenge, parameter-efficient fine-tuning (PEFT) methods [9] have garnered significant attention, focusing on updating only a small fraction of parameters, like adapter [7, 9, 19], prompt tuning [13, 21, 26]. Among these methods, Low-Rank Adaptation (LoRA) [10] has emerged as a groundbreaking technique for LLMs. LoRA introduces a parallel low-rank adapter to the weights of linear layers as shown in Fig. 1(b), reducing memory overhead and computational costs during fine-tuning. However, there are still limitations to LoRA, including activation memory consumption and a performance gap compared to FFT. These limitations have led to the development of many LoRA variants [2, 3, 4, 6, 17, 22, 23, 27, 29, 31, 32].
@@
In this work, we introduce Standard Basis LoRA (SBoRA), a novel approach that further reduces the computational and memory requirements while enhancing the learning capacity. SBoRA achieves this by selectively updating specific rows or columns of the pre-trained weight matrix $\mathbf{W}_{0}$, preserving most of the base model’s weights. The resulting update matrix $\mathrm{\Delta}\mathbf{W}=\mathbf{BA}$ mainly consists of zero rows or columns, indicating that most of the fine-tuned model’s weights $(\mathbf{W}^{\prime}=\mathbf{W}_{0}+\mathrm{\Delta}\mathbf{W})$ remain identical to the pre-trained weights. This localized learning process mirrors the modular organization of the human brain, where specific cognitive functions are localized in distinct regions [5]. For example, the hippocampus is responsible for episodic memory [30]. This design potentially improves knowledge retention and adaptation efficiency.
Let’s assume that the pre-trained weight matrix $\mathbf{W}_{0}$, the projection-down matrix $\mathbf{A}$, and the projection-up matrix $\mathbf{B}$ are given. SBoRA adopts a unique approach, utilizing orthogonal standard basis vectors (one-hot vectors) to construct projection matrices $\mathbf{A}_{sb}$ (for SBoRA-FA) or $\mathbf{B}_{sb}$ for (SBoRA-FB), as shown in Figures 1(c) and 1(d). This approach eliminates the need to store coefficients for one projection matrix, reducing memory usage by approximately 50%. In SBoRA-FA, only matrix $\mathbf{B}$ is updated during fine-tuning, with $\mathbf{A}_{sb}$ and $\mathbf{W}_{0}$ remaining frozen. The weight changes ($\mathrm{\Delta}\mathbf{W}=\mathbf{BA}_{sb}$) reside in a low-rank subspace spanned by $\mathbf{A}_{sb}$, and the projection-down transformation can be efficiently achieved by selecting a subset of input vector samples. SBoRA-FB operates similarly, updating only matrix $\mathbf{A}$ with a pre-defined $\mathbf{B}_{sb}$. Both variants align with LoRA, reducing computational overhead during fine-tuning without introducing extra inference latency.
@@
The primary contributions of this work can be summarized as follows:
  • Introduces SBoRA as a novel PEFT approach for LLMs, which reduces the number of trainable parameters or doubles the rank while improving performance.
  • Presents SBoRA-FA and SBoRA-FB, which use standard basis vectors for sparse updates, keeping most weights unchanged from the pre-trained model.
  • Demonstrates that SBoRA-FA outperforms methods like LoRA and OA in reasoning tasks with half the trainable parameters, and matches DoRA’s performance.
  • Demonstrates effectiveness with quantized LLaMA models, showcasing versatility across different model types.
@@ 2 Related Work
2.1 Full Fine-Tuning (FFT) and Parameter-Efficient Fine-Tuning (PEFT)
Fine-tuning is a fundamental concept in deep learning that enables leveraging pre-trained models for new tasks. However, FFT has limitations, including high storage and computational costs. To address these limitations, PEFT techniques [9] have been proposed, which adapt only a small subset of a model’s parameters to a new task. Early approaches include the use of adapter layers, which add small, trainable modules to each Transformer layer. Another approach is prefix tuning [16], which optimizes a few continuous "prefix" vectors prepended to the input sequence. Both of these approaches significantly reduce the memory and computation needed to fine-tune large models. However, they also have their limitations, such as extra inference latency in the case of adapter layers, and difficulty in optimization for prefix tuning.
@@ 2.2 Low-Rank Adaptation(LoRA)
LoRA [10] is a prominent PEFT method that approximates full-rank updates with low-rank updates within the FFT domain. Given a pre-trained parameter matrix $\mathbf{W}_{0}\in\mathbb{R}^{d\times r}$, LoRA uses two low-rank matrices, $\mathbf{A}\in\mathbb{R}^{r\times k}$ and $\mathbf{B}\in\mathbb{R}^{d\times r}$, to compute the weight update $\mathrm{\Delta}\mathbf{W}\in\mathbb{R}^{d\times k}$. The output vector $\mathbf{h}\in\mathbb{R}^{d\times 1}$ of a LoRA linear layer can be expressed as:
@@
LoRA ensures $\mathrm{\Delta}\mathbf{W}$ is initialized to zero by initializing $\mathbf{A}$ with a uniform distribution and $\mathbf{B}$ with zero. The low-rank decomposition of $\mathrm{\Delta}\mathbf{W}$ into $\mathbf{BA}$ reduces the rank of $\mathrm{\Delta}\mathbf{W}$, making it more efficient than full-rank updating in FFT. LoRA’s low-rank updating approach has demonstrated comparable performance to full-rank updating in tasks like text classification and instruction tuning, while reducing memory usage and computational requirements. It also simplifies deployment in multi-task scenarios and can match or exceed FFT performance using only a fraction of the total parameters.
@@ 3.2 Regional Weight Update
An interesting property of the proposed SBoRA is that the merged weight matrix $\mathbf{W}^{\prime}$ of the fine-tuned model is only regionally updated, with most of the weights remaining unchanged from the original pre-trained weight $\mathbf{W}_{0}$ as depicted in Fig. 2. Specifically, the updated weight matrix $\mathbf{W}^{\prime}$ of SBoRA-FA can be formally expressed as:
@@ 4 Complexity Analysis of SBoRA
SBoRA can be implemented flexibility, with a proposed sampling-based multiplication (sampleMul) for improved computation and memory efficiency. While SBoRA can be implemented similarly to LoRA using matrix multiplication, it offers optimizations by representing the standard basis matrix $\mathbf{A}_{sb}$/$\mathbf{B}_{sb}$ as a $r$-dimensional vector with integers corresponding to the indices of the selected subspace basis vectors. These indices are initialized by randomly selecting $r$ unique numbers from 1 to $k$ for SBoRA-FA or from 1 to $d$ for SBoRA-FB. This implementation achieves approximately 50% reduction in parameter storage and around half of the gradient storage due to the frozen standard basis matrix. Table 1 compares the memory requirements of SBoRA and LoRA.
@@
SBoRA offers potential computation speedups through efficient implementation of matrix operations which leverages one-hot standard basis vectors. For SBoRA-FA, the multiplication $\mathbf{A}_{sb}\mathbf{x}$ can be efficiently implemented as a sampling operation ($\mathbf{x}[seq,\mathbf{A}_{sb}]$), which selects specific elements from $\mathbf{x}$ based on the indices in $\mathbf{A}_{sb}$. The forward process of SBoRA-FA with sampleMul can be expressed as:
@@
This sampling based multiplication (sampleMul) can be more efficient than standard matrix multiplication, especially for larger rank values. Similarly, for SBoRA-FB, the projection-up transformation can be implemented as a scatter operation using $index\_add$($\cdot$). The forward process of SBoRA-FB with sampleMul can be expressed as:
@@
As shown in the equation, projection-down vector $\mathbf{z}$ is directly added into the pre-trained weight optput through index addition method.
These optimizations reduce computation complexity by saving one matrix multiplication operation compared to LoRA. During this process, two matrix multiplication operations $\mathbf{B}(\mathbf{A}_{sb}\mathbf{x})$ / $\mathbf{Ax}$ and $\mathbf{W}_{0}\mathbf{x}$, and one matrix addition operation of $\mathbf{W}_{0}\mathbf{x}+\mathbf{BA}_{sb}\mathbf{x}$ / $\mathbf{W}_{0}\mathbf{x}+\mathbf{B}_{sb}\mathbf{Ax}$ are involved. Table 2 provides an analysis of computational requirements for SBoRA using sampleMul compared to LoRA, assuming input $\mathbf{x}$ in the shape of $[d\times 1]$.
Table 2: Computational analysis and comparison of LoRA and SBoRA’s sampleMul implementation, with a pre-trained weight shape of $k\times d$ and a LoRA rank of $r$.
@@ 5 Experiment
We conducted a comprehensive evaluation of SBoRA’s effectiveness across a range of tasks. To begin with, we conducted a comparative analysis by fine-tuning LLaMA models [25] of different versions and scales on both commonsense reasoning tasks and arithmetic reasoning tasks to assess the performance of SBoRA-FA and SBoRA-FB in comparison to other established PEFT methods. Furthermore, recognizing the growing interest in quantized models, we also evaluated SBoRA on the quantized versions of LLaMA models, utilizing the techniques introduced in QLoRA [4].
@@ 5.1 Evaluating SBoRA on Commonsense Reasoning Tasks
We evaluated SBoRA-FA/FB against LoRA, DoRA, Orthogonal Adaptation (OA), and LoRA-FA approach on the LLaMA-7B/LLaMA3-8B language models for commonsense reasoning tasks. We followed the evaluation method introduced in DoRA [17] and LLM-Adapter [11] which used a comprehensive training dataset created by aggregating eight tasks, evaluating models on individual testing datasets for each task. The results of GPT-3.5 are included for reference as shown in Table 3.
For a comprehensive comparison, we initially fine-tuned models with SBoRA-FA/FB, OA, and LoRA-FA, both using a double rank of 64 to maintain a nearly identical number of trainable parameters (TP) as LoRA and DoRA (rank=32), with only additional storage for standard basis indices. We then conducted experiments with rank 32 for SBoRA-FA/FB, OA, and LoRA-FA, and rank 64 for other baselines to investigate if SBoRA and OA could achieve better performance with half the trainable parameters. Consistency was maintained by using one training epoch and learning rate of 2e-4 for all experiments.
Table 3: Comparative analysis of LLaMA-7B/LLaMA3-8B with different Parameter-Efficient Fine-Tuning methods, evaluating on the commonsense reasoning task. Results of GPT-3.5 are provided in the first row for reference. We report accuracy (%) for eight sub-tasks and average accuracy (%), with higher values indicating better performance. Column headers denote TP for trainable parameters and $r$ for rank.
@@
In LLaMA-7B, SBoRA-FA (rank=64) outperforms all baseline methods with a similar number of trainable parameters, including LoRA (rank=32), DoRA (rank=32), OA (rank=64) and LoRA-FA (rank=64). SBoRA-FA with rank 32 demonstrates commendable performance by still outperforming OA (rank=32) and LoRA-FA (rank=32), while achieving the same accuracy as DoRA (rank=32). It lags behind the highest accuracy achieved by LoRA (rank=32) by only 1.1%. Despite having approximately half the trainable parameters, SBoRA performs comparably to models with the same rank. In LLaMA3-8B, the top two results are achieved by SBoRA-FA with rank 64 and 32, respectively.